OpenStudy (anonymous):

Sabrina is fencing her yard, which is shaped as a rhombus. She used (6x-7) meters of fencing material for one side and (3x+5) meters for the adjacent side. what is the perimeter of the yard?

OpenStudy (p0sitr0n):

6x-7 = 3x+5, as rhombus has 4 equal sides 6x-3x = 5+7 3x = 12 x =4m 4m is one side, P = 4 sides, so 4 x 4m = 16m